OpenStudy (darkigloo):

An 8 Newton block is accelerating down a frictionless ramp inclined at 15 degrees. What is the magnitude of the net force causing the block's acceleration? The answer is 2.1N but I don't know why.

OpenStudy (sidsiddhartha):

what are u getting?

OpenStudy (sidsiddhartha):

 the force should be F=mg*sin(theta)=8*9.81*sin15=20.31